OpenStudy (anonymous):

giving medal for answer ,helppp? What are the possible dimensions for a triangle with an area of 32 cm^2? a. b = 4 cm h = 8 cm c. b = 8 cm h = 8 cm b. b = 2 cm h = 16 cm d. b = 6 cm h = 8 cm

OpenStudy (lncognlto):

A of a triangle = 1/2 b x h, so which one fits?

OpenStudy (anonymous):

(/.\) idk :x

OpenStudy (anonymous):

area A=1/2*b*h b*h=2A

OpenStudy (lncognlto):

So b x h = 64. so b and h are = ?

OpenStudy (anonymous):

c.?

OpenStudy (lncognlto):

Yep thats correct! :)

OpenStudy (anonymous):

ok thankss